A principal at a certain school claims that the students in his school are above average intelligence. A random of thirty (30) students IQ scores have a score of 112. Is there sufficient evidence to support the principal’s claim?

The mean population IQ is μ= 100 with a standard deviation of 15.

Alpha = 0.01

Based on the above data we should

 

       a. not reject H0.            b. reject H0.




1
Expert's answer
2021-12-06T16:11:48-0500

The hypotheses tested are,

"H_0:\mu=100\space vs \space H_A:\mu\gt 100"

We are given that,

"\bar{x}=112,\space n=30,\space \sigma=15"

The test statistic is given as,

"Z_c=(\bar{x}-\mu)/(\sigma/\sqrt{n})"

"Z_c=(112-100)/(15/\sqrt{30})=12/2.7386=4.3818"

"Z_c" is compared with the table value at "\alpha=0.01" given as, "Z_{\alpha}=Z_{0.01}=2.33".

The null hypothesis is rejected if "Z_c\gt Z_{0.01}."

Since "Z_c=4.3818\gt Z_{0.01}=2.33," we reject the null hypothesis and conclude that there is sufficient evidence to indicate that the students in his school are above average intelligence at 1% level of significance.


The correct answer is b. reject H0